OpenStudy (anonymous):

Suppose that an object is 100 pounds when it is at sea level. Find the value of C that makes the equation true. (Sea level is 3,963 miles from the center of the Earth.)

OpenStudy (turingtest):

the equation you say?

OpenStudy (anonymous):

yes

OpenStudy (turingtest):

:/ It would help to know the equation...

OpenStudy (anonymous):

my bad w=Cr^2

OpenStudy (turingtest):

since we are trying to find C we need to solve this equation for C. What have you got for that?

OpenStudy (anonymous):

the first part of the question ask me to solve for r i did so do i need to solve for c

OpenStudy (anonymous):

c=r^2w

OpenStudy (turingtest):

Well I don;t know about the last problem, but here you need C, so yes solve for C. Was the original equaion \[w=C/r^2\] or \[w=Cr^2\]???

OpenStudy (anonymous):

the first one

OpenStudy (anonymous):

no its w=Cr^-2

OpenStudy (turingtest):

Thought, so. Same thing. In any case you are right, solving for C gives\[C=w r^2\]now plug in the numbers and you get your answer.

OpenStudy (anonymous):

what number

OpenStudy (turingtest):

what does w represent? what does r represent? the numbers are given in the question, just plug them into the equations.

OpenStudy (turingtest):

*equation

OpenStudy (anonymous):

Hint: Pay attention to the units of measure. You may have to convert from feet to miles several times in this assignment. You can use 1 mile = 5,280 feet for your conversions. 1. Many people know that the weight of an object varies on different planets, but did you know that the weight of an object on Earth also varies according to the elevation of the object? In particular, the weight of an object follows this equation: , where C is a constant, and r is the distance that the object is from the center of Earth. a. Solve the equation w= Cr^2 for r. r = √wC/w b. Suppose that an object is 100 pounds when it is at sea level. Find the value of C that makes the equation true. (Sea level is 3,963 miles from the center of the Earth.) c. Use the value of C you found in the previous question to determine how much the object would weigh in i. Death Valley (282 feet below sea level). ii. the top of Mount McKinley (20,320 feet above sea level).

OpenStudy (anonymous):

Now can some one please help me out Iam really trying

OpenStudy (turingtest):

why have you not plugged in the numbers to get the answer for part b, I set you up for it?

OpenStudy (anonymous):

im trying something is not right

OpenStudy (anonymous):

how do i plug them together

OpenStudy (turingtest):

\[C=w r^2\]what is e? what is r? plug those in. Do you not know the meaning of of plugging in? It simply means put the number for w where the w is and the number for r where the r is Show me what you are doing please, so I can see your error.

OpenStudy (anonymous):

ok

OpenStudy (turingtest):

what is w* it should be above

OpenStudy (anonymous):

w=c/r^2

OpenStudy (turingtest):

no you just rewrote the original problem. Plugging in is putting numbers for a given variable in place of that variable in the equation. For example in the equation y=x-7 if we PLUG IN x=10 we get y=10-7-3 so apply the same to your equation C=wr^2 with w and r, and tell me what you get.

OpenStudy (turingtest):

*sorry y=10-7=3

OpenStudy (anonymous):

oh ok whendo ill let u knoww what i came up with and thank you for all ur help if i can give u a nother medal i would

OpenStudy (turingtest):

do you not yet have the answer? let me give this one away: the weight is w=100 distance at sea level is r=3963 plugging this into our formula gives\[C=w r^2=100(3963)^2=?\]you can finish from here?

OpenStudy (anonymous):

i get no solution

OpenStudy (anonymous):

my bad give me a few min working on it

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

i got c=wr^2 =1570536900 is that correct

OpenStudy (anonymous):

C=wr^2= 1570536900 is that right so far? whats my next step?